A 3% solution of sulfuric acid was mixed with an 18% solution to produce an 8% solution. how much 3 % solution and 18% solution were used to produce 15 liters of 8% solution?
Answer by checkley77(12844)

.18X+.03(15-X)=.08*15
.18X+.45-.03X=1.2
.15X=1.2-.45
.15X=.75
X=.75/.15
X=5 lITERS OF 18% SOLUTION IS USED.
15-5=10 lITERS OF 3% SOLUTION IS USED.
PROOF:
.18*5+.03*10=.08*15
.9+.3=1.2
1.2=1.2